One integer is seven less than another integer. The product of these two integers is 228. What are the two intergers?
1
Expert's answer
2019-04-09T10:36:58-0400

x - the first integer

(x-7) - the secend integer

we have equation:

"x*(x-7)=228"

x2-7*x-228=0

D=(-7)2-4*1*(-228)=961

x1=(7-31)/2=-12

x2=(7+31)/2=19

the secend integer x-7=-12-7=-19 or x-7=19-7=12

answer -12; -19; or 12;19